Find Kp for the following reaction. SO2Cl2--------> SO2 + Cl2 if sulphuryl chloride decomposes to the extent of 91.2% at 102°C and total pressure 1 atm. Ans. Kp = 4.94 atm

Grade:12th pass

1 Answers

Md Mushahid
4 years ago
Find Kp for the following reaction. SO2Cl2--------> SO2 + Cl2 if sulphuryl chloride decomposes to the extent of 91.2% at 102°C and total pressure 1 atm. Ans. Kp = 4.94 atm
Ans :- 
k_{p} =k_{c}\left ( RT \right )^{\Delta ng}
SO_{2}Cl_{2} \rightleftharpoons SO_{2}+Cl_{2}
\Delta ng=2-1=1
k_{p} =k_{c}\left ( RT \right )^{1}
\frac{^{k_{p}}}{k_{c}} =RT
3 =RT
T =\frac{3}{R}